November 1997  ICTCM X (Chicago)  8

Nonlinear Boundary Conditions

> ODE := diff( y(x), x$2 ) + y(x) = 0:
> BCnl := y(0)=2, y(Pi/2)^2 + D(y)(Pi)=1;

BCnl := y(0) = 2, y(\frac{1}{2} π)^{2} + D(y)(π) = 1

> dsolve( { ODE, BCnl }, y(x) );
> BVPsolve( { ODE, BCnl }, y(x) );

y(x) = 2 cos(x) + RootOf( Z^{2} − Z − 1) sin(x)